### v2

# Foundational Automatic Evaluators: Scaling Multi-Task Generative Evaluator Training for Reasoning-Centric Domains

Abstract
Finetuning specialized generative evaluators has emerged as a popular paradigm to meet the increasing demand for scalable evaluation during both training and test-time. However, recent work has largely focused on applying new methodology, such as reinforcement learning (RL), to training evaluators, shying away from large-scale, data-driven development. In this work, we focus on data scaling, curating a set of 2.5M samples spanning five unique evaluation tasks (pairwise, step-level, reference-free and reference-based verification, and single rating) and multiple domains focused on reasoning evaluation. With our data, we train Foundational Automatic Reasoning Evaluators (FARE), a family of 8B and 20B (with 3.6B active) parameter evaluators, with a simple iterative rejection-sampling supervised finetuning (SFT) approach. FARE-8B challenges larger specialized RL-trained evaluators and FARE-20B sets the new standard for open-source evaluators, surpassing specialized 70B+ evaluators. Beyond static benchmarks, we evaluate FARE in real-world tasks: As inference-time rerankers, FARE-20B achieves near-oracle performance on MATH. As verifiers in RL training, FARE improves the downstream RL-trained model performance by up to 14.1% vs. string-matching verifiers. When initialized from FARE, a continually-finetuned FARE-Code outperforms gpt-oss-20B by 65% on evaluating test-case quality.
[Uncaptioned image]   The FARE family of evaluators
Figure 1: Overview of our work. We curate 2.5M multi-task, multi-domain training samples (left) and use large-scale iterative rejection sampling SFT to train FARE, a family of automatic evaluators (top right). We evaluate FARE on static benchmarks and on various real-world downstream tasks.
1 Introduction
The past two years has seen the rapid adoption of large language models (LLMs) as automatic evaluators in response to demands for scalable evaluation of LLM outputs. LLM-based evaluators serve as judges for popular benchmarks (dubois2024length; zheng2023judging), generative reward models for preference optimization (yuan2024self; wu2024meta), and verifiers/critics in inference-time scaling settings (mcaleese2024llm; zhou2025evaluating). The widespread integration of evaluators into nearly every phase of the LLM development cycle (wu2025sailing) demands evaluators that can handle multiple evaluation tasks while operating effectively across diverse domains.
Different settings require different evaluation abilities: Alignment needs evaluators capable of comparing different responses, i.e., pairwise evaluation, whereas monitoring model outputs requires finding minute mistakes, i.e., step-level evaluation. More recently, generative evaluators are tasked with providing reward signals during reinforcement learning (RL) training (team2025kimi; jiang2025pag), and are expected to grow in importance as RL moves towards unverifiable domains (gunjal2025rubrics; jayalath2025compute) in complex reasoning settings (ke2025survey; ferrag2025llm). As evaluators take central roles in training and evaluating the next generation of models, they must be flexible enough to evaluate as the setting demands.
Compounding the challenges of multi-task evaluation is the expanding number of domains requiring evaluation: RL-training has quickly moved from math reasoning (e.g., yu2025dapo) to general-purpose reasoning (ma2025general) (e.g., history or economics). Agentic settings introduce additional wrinkles: With autonomously acting single agents (openai2025deepresearch; nguyen2025sfr; wei2025webagent) and complex multi-agent workflows (openmanus2025; alzubi2025opendeepsearch) now being set free to browse the web and act on behalf of users with minimal oversight, evaluators must assess not only agent reasoning, but also proposed tool-use. These systems, sometimes built with intricate model-generated (hu2024automated; zhang2024aflow; ke2025mas) interdependencies, are bottlenecked, in part, by subpar evaluation (cemri2025multi).
Unfortunately, recent work in the open-source automatic evaluation community has failed to meet these twin demands of multi-task, multi-domain evaluators, opting instead in training task-specialized evaluators at relatively small data scales. We break this trend by scaling up data, curating 2.5M multi-task, multi-domain training samples that emphasize reasoning settings. As shown in Fig. 1, our data mix covers 5 distinct tasks and various domains like math, code, tool-use evaluation, and natural language reasoning. With our data, we train, Foundational Automatic Reasoning Evaluators (FARE), two best-in-class evaluators. As shown in Fig. 1, our contributions are
  • Multi-task, multi-domain dataset: We curate a large-scale, multi-task training set with an emphasis on reasoning-centric settings. We supplement existing human-and model-annotated data with synthetic data created from challenging new seed datasets.
  • Scalable learning via iterative rejection sampling: We show that iterative rejection sampling supervised finetuning (RS-SFT) is a stable approach for training evaluators at scale. The semi-online nature of RS-SFT avoids problematic teacher model distribution shifts while bringing computationally stable and efficient model updates. Through ablations, we quantify the impact of training pipeline features like quantity of direct judgment data and the use of a continuous curriculum.
  • The FARE family of evaluators: We train FARE-8B and FARE-20B and rigorously assess them with 7 challenging benchmarks and 3 practical downstream settings: test-time response reranking, RL-training verification, and domain-specific continual finetuning.
Our trained models are both well-rounded and high-performing. Out of the box, FARE improve generator performance at test-time, achieving near oracle reranking performance on MATH, and provide clear rewards during general-domain RL training, boosting downstream performance by 14.1% over typical string-matching verifiers. With minimal continual training, FARE can be adapted to specific domains like code, beating gpt-oss-20B by 65% in code test-case quality evaluation.
2 Background and Related Work
An automatic evaluator (AE) $\pi_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ maps input $x=(p,q,\mathcal{R})\in\mathcal{X}$ to output $y=(c,j)\in\mathcal{Y}$. Input $x$ consists of $p$, the evaluation protocol that specifies both the evaluation task (e.g., pairwise comparison, verification) and evaluation rubric, $q$, the original question, and $\mathcal{R}$, set of model responses to be evaluated. The output $y$ consists of a natural language critique $c$ and final judgment $j$. The AE may also be prompted to omit the critique $c$ and directly output the judgment $j$, which we denote as $y=(\emptyset,j)$. The specific evaluation protocol $p$ determines the elements of set $\mathcal{R}$ and the exact form of judgment $j$. For example, in pairwise evaluation, $\mathcal{R}$ consists of two responses $\{r_{1},r_{2}\}$ and the judgment is a choice between the two (“A” or “B”), whereas in single-rating, $\mathcal{R}$ consists of a single response $r$ and the judgment is an integer on a 1-5 scale. In this work, we focus on training automatic evaluators capable of the 5 evaluation tasks shown in Fig. 1:
  • Pairwise comparisons: Given response set $\mathcal{R}=\{r_{1},r_{2}\}$, the AE selects the better of $r_{1}$ and $r_{2}$.
  • Step-level evaluation: Given response set $\mathcal{R}=\{r_{\text{[steps]}}\}$, where $r_{\text{[steps]}}$ is a single model response broken down into steps, the AE identifies step-level errors.
  • Reference-based verification: Given response set $\mathcal{R}=\{r_{\text{cand}},r_{\text{ref}}\}$, where $r_{\text{cand}}$ is the candidate and $r_{\text{ref}}$ is the reference, the AE determines if $r_{\text{cand}}$ is correct based on $r_{\text{ref}}$.
  • Reference-free verification: Given response set $\mathcal{R}=\{r\}$, the AE determines if $r$ is correct.
  • Single rating: Given response set $\mathcal{R}=\{r\}$, the AE assigns an integer score to $r$.
Past work in generative automatic evaluators. Capable LLMs, like GPT-4, were originally prompted as scalable evaluators (wang2023chatgpt; liu2023g; fu2024gptscore; chiang2023can). Subsequent analysis revealed pitfalls of prompted approaches, like biases with respect to position (wang2023large; li2023generative), length (zeng2023evaluating; park2024offsetbias), or self-preference (panickssery2024llm). Finetuning specialized evaluators emerged as a result, with early approaches using teacher model outputs to do supervised finetuning (SFT) (kim2023prometheus; kim2024prometheus; li2023generative; park2024offsetbias; skyworkcritic2024) or direct preference optimization (DPO) (hu2024themis; ye2024beyond), often focusing only on one or two evaluation tasks. More recent methods moved to reasoning models as teachers (khalifa2025process).
vu2024foundational; wang2024direct; cao2024compassjudger; alexandru2025atla train foundational evaluators at larger data scales with multi-protocol capabilities via offline training methods like SFT or DPO. Such approaches take inspiration from general-purpose, large-scale multi-task learning (sanh2021multitask; raffel2020exploring; wei2021finetuned), which showed broad generalization capabilities emerge with the scaling of training data. Foundational evaluators likewise were empirically shown to generalize to unseen evaluation tasks, prompts, and criteria while being more robust to common biases (vu2024foundational; wang2024direct).
Recent work has focused on methodological advances, either using inference-time scaling (liu2025inference; chan2025j1; zhao2025genprm) or online training like reinforcement learning from verifiable rewards (RLVR) (chen2025judgelrm; chen2025rmr1rewardmodelingreasoning; whitehouse2025j1; xu2025j4r; xiong2025stepwiser) to improve evaluator performance. Because RLVR is computationally demanding with relatively brittle training pipelines (dsr1_guo2025deepseek; yang2025qwen3), recent evaluators are typically trained on a small amount of data for a single task. Our work bridges early work in training foundational evaluators with more recent methodological advancements, demonstrating that a simple semi-online training approach enables stable multi-task training at scale.
Desiderata for a new generation of evaluators. Here, we outline our design philosophy for FARE. Beyond accuracy and robustness, we seek efficiency, as many evaluation settings like inference-time reranking or RL rollout verification demand low latency. In contrast to recent long chain-of-thought (CoT) evaluators (chen2025rmr1rewardmodelingreasoning; khalifa2025process), we select base models with either no or very compact “thinking” CoTs. We also explicitly avoid having the evaluator generate reference answers. Past work has used evaluators to generate references during evaluation (zheng2023judging; li2024crowdsourced) or training rollout (chen2025rmr1rewardmodelingreasoning). Not only does this risk severely degrading performance when the reference is wrong (krumdick2025no), it also converts the relatively easier evaluation task into a harder generation task (zhou2025variation).
3 FARE: data and training recipe
3.1 Data curation
We use two data approaches for curating our final training mix: Using Existing high quality training datasets created for evaluator and preference finetuning and generating Synthetic datasets through programmatic error injection and a generate-then-grade strategy.
Existing data consists of training samples from proven sources that have produced effective evaluators (vu2024foundational; cao2024compassjudger). These datasets consist of high quality annotations from humans and frontier LLMs, and cover evaluation tasks like step-level, single-rating, and pairwise evaluation and domains like chat quality, code, and safety. Following wang2024direct, we largely focus our data collection on modern (2024 and beyond) datasets, as these datasets contain the most up-to-date model responses and fresh annotations. Beyond evaluator-specific training, we take advantage of existing preference fine-tuning datasets used for RLHF (instructgpt_ouyang2022training) and DPO training (dpo_rafailov2023direct), converting these directly into pairwise evaluation samples. In domains with objectively correct answers, e.g., math, we also create verification training data with positive responses as correct/reference responses and negative responses as incorrect responses.
We hand-craft evaluation rubrics for each source dataset that follow annotation instructions given to human annotators or models, if existing. If such original instructions do not exist, we write custom evaluation rubrics for each source dataset based on the data composition and domain. § E.2 provides an example rubric. Existing data lays a solid foundation, with 1.4M samples already dwarfing data scales found in recent work. However, upon inspection, we found three clear shortcomings: (1) Newly relevant tasks, like verification, were underrepresented. (2) Existing pairwise task data focused largely on chat-related topics and less-so on reasoning-relevant domains. (3) Questions and responses from newer, challenging datasets produced to meet the needs of reasoning-focused RL training were absent. To address these limitations, we supplement with synthetic data.
Figure 2: Breakdown of our curated training dataset of 2.5M samples by task (left), domain (center), and curation approach (right). Domain breakdown excludes step-level data, which is entirely math.
Synthetic data is generated from a diverse set of challenging seed datasets using two approaches:
  • Programmatic error injection. We employ programmatic error injection when applicable, such as tool-use and function-calling data. For example, to create pairwise tool-use data, we inject errors (e.g., type error, extra argument, syntax errors) in correct function calls. This approach increases the amount of tool-use evaluation data, adding both pairwise and verification samples.
  • Generate-then-grade. Here, we leverage a mix of recent and established training datasets comprised of question $q$ and verifiable ground-truth answer $a$. We sample up to $20$ responses per $q$ from various generator models, then grade each responses based on $a$. After grouping responses by correctness, we create verification and pairwise samples. We use 12 unique generators from 6 model families, covering reasoning and non-reasoning models. This ensures that FARE are trained on a diverse array of model responses, enabling better generalization across distinct response distributions (singh2025shelf). Generate-then-grade enables us to incorporate problems from recent, challenging datasets covering frontier math and reasoning tasks. This enhances the quality of our pairwise and verification data with difficult-to-evaluate reasoning-focused samples.
Our final dataset comprises 2.5M training samples; an overview is shown in Fig. 2, with breakdowns by task, domain, and curation approach, and exact dataset sources are described in App. B.
3.2 Model training
General training recipe. We aim to train an automatic evaluator $\pi_{\theta}$, which we call the policy model parametrized by $\theta$. Our data curation process yields a training dataset of $N$ samples $\mathcal{D}=\{(x_{i},j_{i}^{\star})\}$, where $j_{i}^{\star}$ denotes the ground-truth judgment. Corresponding ground truth critiques $c^{\star}$ are not typically available in evaluator training data. As a result, past work has resorted to offline teacher-model based approaches or online RL-based approaches, as discussed in § 2.
These established approaches have their limitations: Teacher models introduce distribution shifts with respect to the policy model, a common problem with imitation learning (ross2011reduction), with choice of teacher model having a large impact on downstream performance (guha2025openthoughts). On the other hand, RL training is compute and time-intensive, making it difficult to scale to large data quantities, with past work (liu2025prorl) only exploring small (1.5B parameter) models via ad-hoc interventions like reference policy resetting.
We borrow desirable qualities from both paradigms and use semi-online iterative rejection sampling SFT (RS-SFT) (touvron2023llama; dong2023raft), which was recently shown to be competitive with RLVR approaches (xiong2025minimalist). RS-SFT avoids sub-optimal distribution shift by finetuning on correct evaluation traces produced by the policy model while employing a computationally lightweight policy update step. This enables simple yet stable scaling to millions of training samples without sampling from a teacher model. An overview of our approach is shown in Fig. 1.
Concretely, we train $\pi_{\theta}$ as follows. We split our $N$ training samples into disjoint rollout batches $\mathcal{B}_{t}=\{(x_{i,t},j_{i,t}^{\star})\}$ of fixed size $N_{\text{rollout}}$ and initialize the initial policy $\pi_{\theta_{0}}$ to be an existing post-trained LLM (e.g., gpt-oss-20B). Then for step $t=0,\ldots,T-1$, we perform the following:
  • Rollout from previous policy: For inputs $x_{i,t}$ from $\mathcal{B}_{t}$, sample $K$ responses per input from policy $\pi_{\theta_{t}}$, denoted $\{\hat{y}_{i,t}^{(1)},\ldots,\hat{y}_{i,t}^{(K)}\}$.
  • Rejection sampling: For each of the $K$ responses $\{\hat{y}_{i,t}^{(1)},\ldots,\hat{y}_{i,t}^{(K)}\}$, we determine correctness with ground-truth judgment $j_{i,t}^{\star}$. For inputs with correct responses, one randomly chosen response is kept. Any inputs without correct responses are discarded. We denote the collected set of inputs and corresponding correct responses as $\mathcal{D}_{t}$.
  • Policy update: Use $\mathcal{D}_{t}$ to update the policy weights via SFT, initializing with $\theta_{t}$:
    $\theta_{t+1}=\operatorname*{arg\,max}_{\theta}\sum\limits_{(x,y)\in\mathcal{D}_{t}}\log\pi_{\theta}(y|x)$ (1)
Our approach draws inspiration from algorithms such as STaR (zelikman2022star) and RAFT (dong2023raft), with some key differences. STaR notably re-initializes training from $\pi_{\theta_{0}}$ for each iteration $t$ and samples only one greedy response per input prompt, while RAFT relies on an external reward model to rank generated outputs. Because the automatic evaluation setting is inherently verifiable, we omit the need for a reward model to rank sampled responses.
Specific to evaluators, the Self-Taught Evaluator (STE) paradigm of wang2024self and follow-up EvalPlanner (saha2025learning) are closely related to our approach. STE follows STaR with policy re-initialization and EvalPlanner uses multiple SFT and DPO training runs per iteration $t$. Further, these works use in-the-loop synthetic data generation, sampling responses to a small number (<25<25K) of seed questions from a fixed generator. Pairwise samples are then created from correct/incorrect responses and used to train the model. This data generation approach cannot be adapted to create data for other tasks, like step-level evaluation, fundamentally limiting the task abilities of STE. A secondary concern is a lack of exposure to diverse response distributions; Evaluations in wang2024direct show that scaling training data with a simpler training recipe leads to better generalization across benchmarks compared to STE.
Batch composition. For each rollout batch $\mathcal{B}_{t}$, we select unseen training samples from our curated training dataset, ensuring the task mixture is consistent with global task composition. For example, 33% of our overall training data are pairwise tasks (Fig. 2), so 33% of the input prompts in $\mathcal{B}_{t}$ are sampled from unseen pairwise samples. We then sample $K=4$ responses per sample with a temperature of $0.9$, and determine correctness based on final judgment.
Inclusion of direct judgment data. Past work (wang2024direct; cao2024compassjudger) has showed the importance of including direct judgment data samples to isolate judgment training signal. These are samples where the critique $c$ is omitted, and the input protocol $p$ is modified to prompt directly for a judgment. To precisely control the fraction of direct judgment data, we convert a fixed percent of $\mathcal{D}_{t}$ to direct judgment data by dropping generated critiques and modifying the input prompt accordingly. In App. D, we ablate the proportion of direct judgment data and show such data enables FARE to be prompted to exclude critiques for faster inference.
Per-batch continuous curriculum learning. We additionally use a continuous curriculum in training: For each $(x,y)\in\mathcal{D}_{t}$, we compute the pass percentage from the $K=4$ rollout generations for $x$, then sort the dataset in descending order of pass percentage. That is, samples where all $4$ sampled outputs are correct are used to update the model first, and samples where only $1$ of $4$ sampled outputs are correct are used to update the model last. We find this has negligible impact on pairwise domains but large impacts in step-level evaluation, as we show in App. D.
Base models. We train two models starting from Qwen3-8B-Base (yang2025qwen3) and gpt-oss-20B (agarwal2025gpt), denoted FARE-8B and FARE-20B. We find Qwen3-8B (post-trained) to be over-trained, and therefore cold-start Qwen3-8B-Base from SFT data from Qwen2.5-32B-Instruct, which we denote Qwen3-8B-ColdStart. See § B.2 for additional details.
Table 1: Pairwise evaluation results, with best and second-best performance in each section marked. FARE achieve best-in-class performance, even outperforming frontier models in tool-use evaluation. $\dagger$ indicates that benchmark uses consistent accuracy (25% random baseline).
[…]
Table 2: ProcessBench results, with best and second-best performance in each section marked. FARE-20B almost matches GPT-5 with the same prompt, achieving best-in-class performance, while FARE-8B beats comparably sized generative (Gen.) evaluators.
[…]
4 Experiments
We evaluate FARE on both core benchmarks, static benchmarks for automatic evaluators, and in downstream settings, which simulate real applications of evaluators. We provide descriptions of benchmarks and baselines in App. C and additional ablations and analysis in App. D.
4.1 Core benchmarks
Setup. We evaluate along five diverse aspects: (ireasoning with JudgeBench (tan2024judgebench), ReasoningJudgeBench (RJB) (xu2025j4r), and PPE Correctness (frick2024evaluate), (iibias and robustness with RM-Bench (liu2024rm), (iiitool-use with When2Call (ross2025when2call), (ivstep-level error identification with ProcessBench (zheng2024processbench), and (vreference-based verification with VerifyBench (yan2025verifybench).
For RM-Bench and PPE Correctness pairwise benchmarks, we adopt default evaluation setups, running each benchmark once with a fixed random ordering of responses. For other pairwise benchmarks, we report consistent-accuracy (tan2024judgebench), where each test sample is run twice, swapping the order of response A and response B. If the evaluator selects a different response between runs (i.e., is positionally biased), then the sample is marked incorrect; if the evaluator is consistent, the judgment is graded against the ground-truth. For ProcessBench and VerifyBench, we report F1-score [^1] ProcessBench defines their reported F1-score differently from the traditional F1-score; See this link. and accuracy, respectively. We compare FARE against other finetuned generative and prompted evaluators. We report official numbers from past benchmarks, reporting sources in App. C. If necessary, we run each baseline using its own prompt template. For ProcessBench, we additionally compare against non-generative process reward models (PRMs).
Results. Tables 1, 2 and 3 present our results on pairwise, step-level, and reference-based verification benchmarks, respectively. Our prompts are provided in App. E.
Table 3: Ref.-based verification, with best and second-best performance marked per section. FARE beat general-purpose verifiers in hard settings.
[…]
Table 1 shows that across diverse pairwise benchmarks, FARE exhibit best-in-class performance, outperforming comparably sized baselines. FARE-8B is the strongest small judge, outperforming recently released RL-trained models like J1-8B and RM-R1-14B by 13.71 and 6.57 absolute points on JudgeBench, respectively. FARE-20B challenges strong judges at 20B parameters, outperforming dense 70B-sized judge models despite having 3.5x fewer total parameters and nearly 20x fewer active parameters. The strong performance of FARE across reasoning benchmarks, which span math to scientific domains to causal reasoning, show that our models excel at discerning between objectively correct and incorrect responses. Beyond reasoning settings, FARE are generally robust to subtle, stylistic biases (RM-Bench) while excelling in tool calling evaluation (When2Call), which is increasingly important as agentic workflows grow in popularity.
FARE also are extremely strong step-level evaluators, as shown by ProcessBench performance in Table 2. FARE-8B is the best small-sized generative critic model, outperforming the recently released StepWiser-7B (xiong2025stepwiser), a RL-trained specialized step-level evaluator, by 1.6 points. FARE-20B outperforms the specialized Qwen2.5-Math-72B-PRM by 6.1 points, even matching GPT-5. Most notably, FARE excel on the two most challenging splits, OlympiadBench and OmniMATH, with FARE-8B beating StepWiser-7B by 5.6 points and FARE-20B beating Qwen2.5-Math-72B-PRM by 7.3 points on average. Overall, FARE are not only capable outcome-level evaluators, but are able to find subtle mistakes that manifest at the step-level.
FARE are also capable verifiers, as shown in Table 3, with both models outperforming all reported baselines on VerifyBench-Hard (yan2025verifybench). In particular, FARE-20B excels beats the next best model, GPT-4o, by 12.5 absolute points. As we demonstrate in § 4.2, when used as verifiers during GRPO settings, FARE bring tangible benefits over typical string-matching verifiers.
Scaling inference-time compute. Recently, sampling parallel judgments and aggregating via majority vote, i.e., self-consistency (wang2022self), has been used to improve evaluator performance. In Table 4, we use self-consistency with 32 responses (SC@32) on PPE, comparing with J1 and DeepSeek-GRM (liu2025inference). DeepSeek-GRM also trains a MetaRM to perform judgment re-ranking at test-time. Across most splits, using SC@32 improves performance, with up to a 3.3 point improvement for FARE-20B on the MMLU-Pro split. Even without SC@32, FARE-8B beats DeepSeek-GRM-27B + MetaRM, with the gap widening with extra compute. Similarly, the gap between FARE-20B and J1-70B grows with extra compute. Interestingly, we see that across both our models, performance on MBPP+ slightly degrades, indicating that SC may not be the optimal way to use compute across all domains. Nonetheless, we observe gains in the aggregate.
Table 4: Scaling inference-time compute for FARE typically brings additional gains in performance: The performance gaps between FARE-8B/DeepSeek-GRM and FARE-20B/J1-70B widens.
PPE MMLU-Pro MATH GPQA MBPP+ IFEval Overall
J1-8B w/ SC@32 65.6 $\rightarrow$ 67.5 70.0 $\rightarrow$ 76.6 53.2 $\rightarrow$ 55.7 53.1 $\rightarrow$ 54.6 54.0 $\rightarrow$ 54.9 59.2 $\rightarrow$ 61.9
J1-70B w/ SC@32 79.0 $\rightarrow$ 79.9 86.0 $\rightarrow$ 88.1 65.9 $\rightarrow$ 66.5 66.0 $\rightarrow$ 66.5 67.3 $\rightarrow$ 67.2 72.8 $\rightarrow$ 73.6
[…]
FARE-8B w/ SC@32 69.3 $\rightarrow$ 70.8 79.7 $\rightarrow$ 80.4 58.4 $\rightarrow$ 58.4 55.7 $\rightarrow$ 54.9 55.9 $\rightarrow$ 56.7 63.8 $\rightarrow$ 64.2
FARE-20B w/ SC@32 80.3 $\rightarrow$ 83.6 94.6 $\rightarrow$ 97.3 68.5 $\rightarrow$ 71.1 59.3 $\rightarrow$ 57.3 69.1 $\rightarrow$ 71.1 74.4 $\rightarrow$ 76.6
4.2 Downstream evaluation
Setup. We apply FARE on 3 downstream tasks: (i) Reward model for inference-time scaling, (ii) verifier for GRPO training, and (iii) initialization for continual finetuning for domain-specific evaluation. We provide detailed explanations of our downstream evaluation settings in § C.2.
Figure 3: Best-of-10 performance for Llama-3.1-8B generator across 4 challenging benchmarks with baseline (red line) and oracle (green line) performance with FARE and SFR-Judge (SFR), Skywork Critic (SC), Compass-Judger (CJ), and Self-Taught Evaluator (STE) as baselines. FARE are the best small ($\leq$14B) and large ($\geq$ 20B) reranking models: FARE-20B achieves near oracle re-ranking performance on MATH, while FARE-8B matches 70B judges.
Reward model for inference-time scaling. We use the standardized setup in JETTS (zhou2025evaluating), which provides a set of 10 outputs from various generators and various benchmarks with corresponding correctness labels. Here, automatic evaluators rerank the responses, and performance is measured as the final performance of the evaluator-selected responses. We select the four most challenging benchmarks used in JETTS: MATH (hendrycks2021measuring), CHAMP (mao2024champ), MBPP+ (evalplus), and BigCodeBench (zhuo2024bigcodebench). We compare against strong judge models benchmarked previously on JETTS: SFR-Judge-8B,70B (wang2024direct), Skywork-Critic-8B,70B (skyworkcritic2024), Self-Taught-Evaluator-70B (wang2024self), and CompassJudger-7B,14B (cao2024compassjudger). We utilize the default pairwise reranking setup and prompt FARE to produce a judgment directly without explanation, i.e., $y=(\emptyset,j)$.
Fig. 3 shows best-of-10 performance with Llama-3.1-8B as generator with baseline greedy (red line) and oracle re-ranking performance (green line). FARE produce best-in-class reranking performance, with FARE-8B roughly matching the performance of larger (70B) judges in math settings and outperforming all similar sized judges in coding domains. FARE-20B excels in math domains, approaching oracle-level reranking performance on MATH, beating SFR-Judge-70B and Skywork-Critic-70B by 14 and 21 absolute points, respectively. Similarly, FARE-20B beats 70B+ judges by large margins in challenging coding domains. As we show in § D.6, FARE improve the performance of other generators and FARE-20B improves significantly over gpt-oss-20B as a reranker.
Verifer for GRPO training. We train with WebInstruct-Verified (ma2025general), a multi-domain reasoning dataset, covering math, chemistry, etc. Verifier impact is measured via the downstream performance of the trained policy model on a fixed evaluation suite of MMLU-Pro (wang2024mmlu), GPQA-Diamond (rein2024gpqa), MATH-500, Minerva-Math (lewkowycz2022solving), OlympiadBench (he2024olympiadbench), and AIME24 (Avg@32). We start from Qwen2.5-7B-Base (yang2024qwen2) with the default reward setup as ma2025general (see § C.2), and we compare against training with the string-matching and trained verifier (General-Verifier) from ma2025general.
Figure 4: Performance of downstream GRPO-trained model with FARE as reference-based verifiers. Moving from string based output matching or weaker verifiers to FARE brings tangible performance gains across both natural-language based (e.g., GPQA) and math settings.
Training with FARE-20B as a verifier improves downstream performance from 34.3 to 45.2, a nearly 11 point absolute gain, with improvements coming uniformly across the six benchmarks, as shown in Fig. 4. Notably, with 77% fewer gradient updates [^2] For the same group size, General-Reasoner-7B is trained for 700 steps with rollout batch size 768, whereas we train for 120 update steps with rollout batch size 1024. , our model was able to slightly beat the performance of General-Reasoner-7B (ma2025general) on several benchmarks: 38.9 vs. 38.8 on GPQA-Diamond, 38.4 vs 37.9 on OlympiadBench, and 13.9 vs. 13.8 on AIME24. This shows that using FARE-20B can significantly improve RL training convergence. Further, using FARE-8B and FARE-20B bring 8.8% and 14.1% relative gains over typically used string matching verifiers and 4.4% and 9.4% over General-Verifier, which has been trained with in-training-distribution verification data. These gains appear for both natural language (MMLU-Pro, GPQA) and math domains, showing that FARE can verify complex outputs across multiple challenging domains.
Figure 5: Continual training of FARE-20B for code evaluation with only 15K samples yields larger gains over gpt-oss-20B/120B.
Initialization for domain-specific continual finetuning. We continually finetune FARE-20B for code evaluation with one round of RS-SFT to produce FARE-20B-Code. We train with only 15K pairwise samples randomly chosen from AceCoder (zeng2025acecoder). For evaluation, we use the recently released CodingJudgeBench (jiang2025codejudgebench), a pairwise benchmark covering code generation, code repair, and test-case quality evaluation tasks. Fig. 5, which reports consistent accuracy across the three splits from gpt-oss-20B, FARE-20B, FARE-20B-Code, and gpt-oss-120B for reference, shows that the first two splits are relatively easy, whereas test case evaluation is extremely difficult: gpt-oss-120B achieves only 32%. On the last task, FARE-20B improves over gpt-oss-20B by 11.67 absolute points, highlighting the benefits of large-scale evaluation training. Specialized continual training on coding tasks brings an additional 10.48 absolute point improvement, with FARE-20B-Code outperforming even gpt-oss-120B on average. In all, FARE can be readily adapted for specific applications with a small amount of domain-specific data.
5 Conclusion
Using a curated multi-task, multi-domain training mix and RS-SFT, we train FARE, a family of high performing and well-rounded automatic evaluators. FARE-8B challenges larger specialized evaluators and FARE-20B sets a new standard for locally hosted evaluators. Our evaluations include 7 challenging benchmarks and 3 practical downstream settings where we show that FARE are (1) effective reward models at inference-time, (2) effective verifiers for GRPO training, and (3) strong initializations for continual, domain-specific finetuning.
Appendix
Appendix A Extended Related Work
Here, we cover recent advances in automatic evaluation that are non-generative. Scalar reward models (RMs) (cobbe2021training), which output a single scalar score for a given question $q$ and response $r$ were popularized originally within the context of reinforcement learning from human feedback (RLHF) (instructgpt_ouyang2022training; stiennon2020learning; ziegler2019fine). Here, pairwise human preferences were used to train a RM, which is then applied during RL-based alignment, e.g., with PPO (schulman2017proximal) or DPO (hosseini2024v). Past work has focused on dataset curation (jiang2023llm; liu2024skywork; liu2025skywork; dong2024rlhf) and experimenting with different loss formulations (liu2024skywork; lou2024uncertainty).
The above class of reward models operate at the outcome level, with the entire response being assigned a single score. Step-level reward models, known as process reward models, attempt to provide denser feedback by assigning each “step” in a response a score. Training PRMs relies on step-level labels from humans (lightman2023let) or models (duan2025efficient) or using Monte Carlo simulation to estimate step-level quality (wang2023math; luo2024improve; xiong2025stepwiser; prmlessons). Both approaches have associated drawbacks: Step-level annotation is rarely scalable, while Monte Carlo simulation requires careful filtering to ensure high quality.
While RMs and PRMs remain popular paradigms, recent approaches have found that generative evaluators can better leverage test-time compute for stronger evaluation performance (zhang2024generative; mahan2024generative; liu2025inference), inspiring recent advances in generative evaluators, as discussed in § 2.
Appendix B Data and Training Details
B.1 Training data
We enumerate our training data sources in Table 5. We took efforts to decontaminate our training sets with N-gram matching approaches, following guha2025openthoughts. For the Synthetic data approach, we use 12 unique generators, covering a mixture of weak and strong models: Ministral-8B (ministral8b), Mistral-Small 24B (mistral24b), Gemma 3 12B (team2025gemma), Qwen2.5 7B, 32B (yang2024qwen2), Qwen-QwQ (qwq32b), Qwen3-30B-A3B (yang2025qwen3), Llama-3.1-8B, Llama-3.3-70B (dubey2024llama), GPT-4o (hurst2024gpt), and GPT-4.1-nano, 4.1 (gpt41) for seed datasets with verifiable answers. To increase diversity we randomly select a prompt template from a preset list for each question in the seed dataset and sample multiple responses at varying temperatures (e.g., 0.0, 0.3, 0.5, 0.7,…). For open-ended datasets, such as tool-use datasets, we enumerated common errors found in tool calling outputs, such as invalid input types, missing arguments, malformed json format, etc., and then programmatically injected errors into ground-truth correct responses.
Table 5: Data sources used to create our training set.
Name Task Domain Curation
Approach
ActPRM duan2025efficient Step-Level Math Existing
Beavertails Preference ji2023beavertails Pairwise Safety Existing
Code Preference Pairs Vezora/Code-Preference-Pairs Pairwise Code Existing
DeepScaleR deepscaler2025 Pairwise, Verification Math Synthetic
Folio han2022folio Pairwise, Verification NL Reasoning Synthetic
FoVer kamoi2025training Step-Level Math Existing
HelpSteer wang2023helpsteer Single rating Chat Existing
HelpSteer2 wang2024helpsteer Single rating Chat Existing
HelpSteer3 wang2025helpsteer3; wang2025helpsteer3pref Pairwise Code, NL Reasoning Existing
HH-RLHF Harmless bai2022training; ganguli2022red Pairwise Safety Existing
LAMP chakrabarty2025can Pairwise, Single rating Chat Existing
MATH hendrycks2021measuring Pairwise, Verification Math Synthetic
MemGPT MemGPT/MemGPT-DPO-Dataset Pairwise, Verification Tool-Use Existing
OffsetBias park2024offsetbias Pairwise Chat Existing
ReClor yu2020reclor Pairwise, Verification NL Reasoning Synthetic
StepDPO lai2024step Pairwise, Verification Math Existing
StrategyQA geva2021did Pairwise, Verification NL Reasoning Synthetic
SWEGym pan2024training Verification Code Existing
SWERank reddy2025swerank Verification Code Existing
SynLogic liu2025synlogic Pairwise, Verification NL Reasoning Synthetic
Tulu-V3-IF DPO data lambert2024t Pairwise Chat Existing
WebDPO WebDPO Pairwise, Verification NL Reasoning Existing
When2Call Preference Pairs ross2025when2call Pairwise Tool-Use Synthetic
XLam-60K liu2024apigen Pairwise, Verification Tool-Use Synthetic
B.2 Training details
We train batch size 128 and a constant learning rate of 1e-6 and choose per-iteration rollout batch sizes of 50,000 and 250,000 for FARE-8B and FARE-20B, respectively. In the latter case, we make the practical trade-off of RS-SFT iterations for training speed, reducing the number of times we need to reset the model for rollouts, etc. We use a modified version of the OpenRLHF framework (hu2024openrlhf) for training.
Qwen3 cold-start SFT. Public discussion from members of the Qwen organization indicate that post-trained versions of Qwen3 are difficult to continually finetune, and recommend starting from base models [^3] See, for example, this Twitter/X post: “…Instruct models after RL will pose difficulty for finetuning, but base models I don’t think so…”. Therefore, we opt to cold-start SFT with one iteration of rejection sampling data collected from Qwen2.5-32B-Instruct. As we show in § D.5, while this cold-start model does not match Qwen3-8B, FARE-8B outperforms the non-thinking Qwen3-8B on many static evaluation benchmarks.
We hypothesize that a relatively short, general-purpose alignment phase prior to evaluation-specific finetuning could further improve performance. While we did not attempt this, we believe this line of experimentation is of interest for future work.
Appendix C Benchmark and Baseline Details
C.1 Core benchmarks
For core benchmarks, we select a set of challenging and contemporary benchmarks for evaluating automatic evaluators:
  • JudgeBench (tan2024judgebench): A pairwise benchmark focused on evaluating LLM-as-judge models in reasoning settings, covering math, code, logical reasoning, and knowledge-based reasoning. Responses are generated using GPT-4o.
  • ReasoningJudgeBench (xu2025j4r): A pairwise benchmark that covers more diverse reasoning settings, such as multi-hop, causal, and domain-specific reasoning. Responses are generated using GPT-4o.
  • PPE Correctness (frick2024evaluate): A pairwise benchmark that covers reasoning and instruction following tasks with objectively correct answers, using seed datasets like MATH hendrycks2021measuring, GPQA (rein2024gpqa), MBPP+ (evalplus), MMLU-Pro (wang2024mmlu), and IFEval (zhou2023instruction). Responses are generated using a variety of weaker models, e.g., Gemma-2-9B.
  • RM-Bench (liu2024rm): A pairwise benchmark that evaluates how robust evaluators are to stylistic biases by evaluating on pairs of responses with subtle yet critical differences.
  • When2Call (ross2025when2call): A pairwise benchmark that covers appropriate selection of tools (or refusals) in response to a user prompt. We use the LLM-as-judge test split, which comprises 300 unique prompts. Each prompt has four candidate answers (refusal, direct response, tool call, follow-up question), of which one response is correct. We form all pairs, yielding 900 total pairwise comparisons.
  • ProcessBench (zheng2024processbench): A step-level benchmark that evaluates the ability to identify step-level errors in mathematical reasoning across easy (GSM8K and MATH) and hard (Omni-Math and OlympiadBench) questions.
  • VerifyBench (yan2025verifybench): A reference-based verification benchmark, comprised of Easy and Hard splits, that evaluates verifier ability to identify equivalent final answers.
For all core benchmarks, we utilize officially reported numbers when available. Otherwise, we run the corresponding baseline ourselves, using any prompt templates released with evaluators.
For pairwise benchmarks, we select our baselines from (1) existing multi-task foundational evaluators, (2) recently released RL-trained judge models, and (3) strong-performing specialized judges:
  • RISE-Judge (yu2025improve): Pairwise judges trained with SFT then DPO to perform pairwise evaluation.
  • Self-Taught Evaluators (wang2024self): A pairwise judge trained with iterative SFT with training data generated in the loop.
  • EvalPlanner (saha2025learning): Pairwise judges trained with iterative SFT and DPO on a small seed dataset, with an emphasis on learning how to plan for evaluation tasks.
  • RM-R1 (chen2025rmr1rewardmodelingreasoning): A family of pairwise judges trained with GRPO, initialized from DeepSeek-distilled Qwen models.
  • J1 (whitehouse2025j1): A pairwise and single-rating judge trained with GRPO.
  • CompassJudger (cao2024compassjudger): A family of foundational evaluators trained with large-scale SFT.
  • Atla Selene (alexandru2025atla): A foundational evaluator trained with large-scale preference optimization.
We run gpt-oss variants with low reasoning, as (1) FARE-8B is trained initialized from gpt-oss-20B-low, and (2) evaluation often demands low-latency, making long CoT undesirable if they can be avoided. For GPT-5, we use the default API settings (medium reasoning). For pairwise benchmarks, we use reported scores from whitehouse2025j1; xu2025j4r; liu2025inference.
For ProcessBench, we use officially reported numbers in zheng2024processbench, which includes SkyworkPRMs (he_2024_16998085), Math Shepherd PRM (wang2023math), ActPRM (duan2025efficient), and Qwen-Math PRMs (prmlessons). We additionally report results from generative baselines: RL Tango (zha2025rl) and StepWiser (xiong2025stepwiser). For VerifyBench, we use reported scores from the original paper (yan2025verifybench).
C.2 Downstream settings
Reward model for inference-time scaling. We compare our models against the following baselines, representing best-in-class performers as reported in JETTS. We utilize reported numbers directly except for CompassJudger, which was not included in the original JETTS evaluation. As such, we run CompassJudger ourselves.
  • SFR-Judge-8B, 70B (wang2024direct): A family of multi-task evaluators. Among the highest performing small and large judges on JETTS.
  • Skywork-Critic-8B, 70B (skyworkcritic2024): Two pairwise-specfic evaluators that do not output explanations. Among the highest performing small and large judges on JETTS.
  • Self-Taught-Evaluator-70B (wang2024self): A strong performing large judge model.
  • CompassJudger-7B, 14B (cao2024compassjudger): As described above.
Verifier during GRPO training. We adopt the settings of ma2025general, which train General-Reasoner, a family of reasoning LLMs of varying model sizes using the WebInstruct-Verified training dataset. In particular, we train with standard GRPO, i.e., without dynamic sampling or clip higher modifications, initializing from Qwen2.5-7B-Base. We the same conditional reward setup as General-Reasoner:
  • If the solution parsing fails, then reward is set to 0.5-0.5.
  • If a solution successfully parsed and is deemed correct by the verifier, it is assigned a reward of $1$ plus a length penalty of:
    $-0.05\times\min\{10,|\texttt{len(ground\_truth) - len(model\_response)}|\}.$
The training framework is based on the verl framework (sheng2024hybridflow). We use rollout batch size 1024, max response length of 4096, group size of 8, a temperature of 1.0, a KL coefficient of 0.001, and a learning rate of 5e-7.
Initialization for domain-specific finetuning. We randomly sample 15,000 pairwise samples from AceCoder (zeng2025acecoder) and perform one round of rejection sampling. We adopt training setup of FARE-20B: a direct judgment ratio of 60% and continuous curriculum. We train for 3 epochs with batch size 256 and cosine decay learning rate peaking at 1e-5. We then evaluate on CodingJudgeBench, reporting consistency accuracy.
Note that CodingJudgeBench reports an unconventional pairwise metric, employing Z-score normalization between the two consistency runs. Their implementation is not publicly available, and their paper lacks concrete implementation details. As such, we resort to consistent accuracy, which is more commonly used in pairwise benchmarks, e.g., (tan2024judgebench; li2023generative; xu2025does; xu2025j4r).
Appendix D Ablations, analysis, and additional results
Figure 6: Pairwise positional robustness emerges as at large data scales, as shown with training FARE-8B and an earlier development run initialized with Qwen2.5-7B-Instruct.
Table 6: Ablation study varying proportion of direct judgment data, use of continuous curriculum (8B model), and ablating strategies for using direct judgment data for gpt-oss (20B model).
% direct
judgment data
Qwen3-8B-ColdStart Models gpt-oss-20B Models
Pairwise ProcessBench Average Pairwise ProcessBench Average
30 61.36 52.10 56.73 72.51 83.64 78.08
40 61.14 58.03 59.59 70.82 82.50 76.66
50 61.56 56.85 59.21 71.97 83.00 77.49
60 62.31 54.73 58.52 72.58 84.40 78.49
70 62.67 56.49 59.58 71.59 84.72 78.16
Curriculum Pairwise ProcessBench Average Keep CoT? Pairwise ProcessBench Average
Yes 61.14 58.03 59.59 Yes 67.80 81.64 74.72
No 61.49 53.43 57.46 No 69.81 82.55 76.18
D.1 Training recipe ablations.
In Table 6, we ablate three components of our training recipe and report the average on our five pairwise benchmarks and ProcessBench. First, we train multiple checkpoints using RS-SFT varying the proportion of direct judgment data from 30% to 70%. Direct judgment data affects FARE-8B and FARE-20B differently: endpoints (30% or 60-70%) show the best performance for FARE-20B, whereas performance peaks at 40% for FARE-8B. As such, we choose 40% and 60% to train FARE-8B and FARE-20B.
We also ablate different strategies for training with direct judgment data specific to gpt-oss, which is trained to output an intermediate CoT before responding. We can either keep this CoT with direct judgment data or remove the intermediate CoT by going directly to the assistant turn. The former more closely mimics the training distribution of gpt-oss, but undermines the intended purpose of direct judgment data of isolating outcome correctness training signal, as discussed in § 3.1. The latter is out-of-training-distribution but more effectively isolates training signal. As seen in Table 6 (bottom right), removing the intermediate CoT leads to gains in both pairwise and step-level settings.
Finally, we measure the impact of the continuous curriculum as compared to a random data shuffling strategy with FARE-8B. As shown in Table 6 (bottom left), the continuous curriculum leads to minimal drops in pairwise performance but large gains in ProcessBench.
Table 7: Performance with and without critiques (and CoT for FARE-20B) for pairwise benchmarks (left) and ProcessBench (right). Directly prompting for a verdict universally improves performance for FARE-8B, but degrades performance for FARE-20B.
JudgeBench RJB
PPE
Correctness
RM
Bench
When2Call Average GSM8K MATH
Olympiad
Bench
Omni
MATH
Average
FARE-8B 55.71 51.05 63.8 79.2 80.33 66.0 68.5 67.7 59.9 58.1 63.5
FARE-8B, no critique 60.00 52.60 64.9 81.8 86.55 69.2 68.5 68.6 59.0 58.5 63.7
FARE-20B 64.29 57.05 74.4 90.5 76.67 72.6 89.8 87.8 80.0 79.9 84.4
FARE-20B, no critique or CoT 62.00 55.23 68.9 85.5 89.11 72.1 79.8 73.2 70.0 70.3 73.3
D.2 How does direct judgment prompting affect performance?
Many settings demand low latency, such as inference-time reranking or evaluating rollouts during RL training. Here, we study how performance varies when FARE are prompted to skip the critique $c$ and directly output a judgment $j$. For FARE-20B, this involves additionally skipping the intermediate CoT, directly outputting from the assistant turn, making out-of-distribution relative to gpt-oss-20B’s original training setup. We see that performance improves for FARE-8B, but degrades for FARE-20B. Such results for FARE-8B are in-line with prior work, which finds that direct judgment-like inference leads to minimal drops in performance. For FARE-20B, we hypothesize that the post-training of gpt-oss-20B instills a strong prior in favor of generating intermediate CoT. Even training with direct judgment data, removing such CoT is detrimental. Nonetheless, performance does not universally degrade, with When2Call performance increasing by nearly 13 points.
D.3 Robustness to pairwise positional bias emerges with data scale.
A known issue in pairwise evaluation is inconsistency (wang2023large), a form of positional bias where the evaluator judgment changes based on the order of responses in the input prompt. During training, we observed that our judges become more consistent as a function of data scale; Fig. 6 shows the progression of pairwise consistency on the five pairwise benchmarks in § 4.1 over the course of training for FARE-8B and an earlier training run which was initialized from Qwen2.5-7B-Instruct. Both models steadily become positionally robust over the course of training, with the weaker Qwen2.5-7B-Instruct showing substantial gains. This reveals that scaling evaluator training data can mitigate common judge biases, complementing mitigation strategies that use data augmentation (saha2025learning), label balancing (cao2024compassjudger; wang2024direct), and RL-based reward or algorithmic methods (whitehouse2025j1; xu2025j4r).
Table 8: Single rating performance, with best and second-best performance in each section marked. Despite being trained with a focus on reasoning settings, FARE perform competitively in single-rating evaluation in chat settings.
FLASK BiGGen Bench Average
GPT-4o-mini 0.630 0.600 0.615
Glider-3.8B 0.615 0.604 0.610
FlowAI-Judge-3.8B 0.400 0.460 0.430
Prometheus-2-7B 0.470 0.500 0.485
Auto-J-13B 0.350 0.300 0.325
Themis-8B 0.540 0.580 0.560
SFR-Judge-8B 0.520 0.590 0.555
SFR-Judge-12B 0.590 0.570 0.580
Atla Selene 8B 0.613 0.584 0.599
LMUnit-8B 0.600 0.645 0.623
FARE-8B 0.611 0.616 0.591
GPT-4o 0.690 0.650 0.670
Prometheus-8x7B 0.540 0.520 0.530
SFR-Judge-70B 0.660 0.650 0.655
LMUnit-70B 0.720 0.677 0.699
FARE-20B 0.649 0.616 0.633
D.4 Single-rating evaluation.
We additionally evaluate FARE on Single Rating tasks with BiGGen-Bench (kim2024biggen) and FLASK (ye2023flask), two chat-centric evaluation datasets with human annotated 1-5 ratings. We measure Pearson correlation with human annotations, and report results in Table 8. Single-rating is widely used evaluation task in reasoning settings, and thus constituting the smallest proportion of our training data, as shown in Fig. 2. Nonetheless, FARE are competitive with chat-focused judge models, with FARE-8B outperforming foundational judge models like SFR-Judge-8B and 12B and FARE-20B approaching the performance of SFR-Judge-70B. We use reported values from baseline papers, including LMUnit (saad2024lmunit), Atla Selene (alexandru2025atla), and SFR-Judge (wang2024direct).
Table 9: Comparison of FARE against their initial models and other popular general-purpose models. indicates some results reported in whitehouse2025j1 or zheng2024processbench.
JudgeBench ReasoningJudgeBench PPE Correctness RM-Bench When2Call Avg. consistency ProcessBench
Qwen3-8B-ColdStart 48.29 40.59 60.5 78.07 59.67 72.55 38.3
Qwen3-8B-non-thinking 52.27 43.56 64.8 79.9 64.78 74.04 56.7
FARE-8B 55.71 51.05 63.8 79.2 80.33 82.28 63.5
gpt-oss-20B (low) 59.43 50.51 71.7 89.9 61.33 77.83 73.9
FARE-20B 64.29 57.05 74.4 90.5 76.67 82.92 84.4
gpt-oss-120B (low) 70.29 58.26 77.8 92.0 70.00 84.09 83.4
Deepseek-R1-671B 68.90 58.53 76.5 88.6 81.00 - -
GPT-4.1 66.29 59.68 78.4 87.8 64.00 85.54 77.8
GPT-4o 50.29 45.25 68.9 80.1 67.44 78.02 61.9
o1-mini 64.20 - 71.3 80.8 - - 87.9
D.5 Comparison against general-purpose models
Here, we compare FARE against general-purpose LLMs, selecting popular reasoning and non-reasoning models. Table 9 shows our results. Our cold-start SFT for Qwen3-8B produces the weakest Qwen3 variant, as expected. However, after undergoing iterative SFT, FARE-8B surpasses Qwen3-8B on multiple benchmarks, improving from 43.56 to 51.05 on ReasoningJudgeBench and 56.7 to 63.5 on ProcessBench. Likewise, we are able to improve gpt-oss-20B across the board, yielding substantial improvements in reasoning, tool-calling, and step-level evaluation. The resulting checkpoint approaches gpt-oss-120B on a number of benchmarks.
Table 10: Full results on JETTS. Numbers in bold indicate that the judge reranking was helpful, i.e., performance is greater than baseline (greedy) performance.
Benchmark Generator Baseline Oracle FARE-8B FARE-20B FARE-20B gpt-oss-20B
Model Performance Performance [CritiquePrompt] [CritiquePrompt]
MATH Llama-3.1-8B-Instruct 24.70 53.47 35.73 50.83 49.85 29.83
Llama-3.1-70B-Instruct 43.81 68.35 53.47 65.41 64.66 51.06
Qwen2.5-32B-Instruct 57.10 78.17 65.03 74.32 74.24 61.56
Qwen2.5-72B-Instruct 62.99 82.78 70.17 79.98 78.70 70.32
GSM8K Llama-3.1-8B-Instruct 85.67 96.44 92.04 94.77 94.77 93.78
Llama-3.1-70B-Instruct 95.53 98.48 96.37 96.74 96.74 96.06
Qwen2.5-32B-Instruct 95.22 98.56 96.21 96.29 96.74 95.75
Qwen2.5-72B-Instruct 95.68 97.88 95.98 95.75 95.98 95.53
CHAMP Llama-3.1-8B-Instruct 29.26 60.00 34.07 44.07 42.22 35.93
Llama-3.1-70B-Instruct 47.41 71.48 51.85 58.52 56.67 55.56
Qwen2.5-32B-Instruct 75.19 85.56 70.00 77.78 79.26 74.81
Qwen2.5-72B-Instruct 71.48 85.56 70.00 73.70 74.81 67.78
MBPP Llama-3.1-8B-Instruct 54.50 76.46 59.79 68.78 68.25 63.49
Llama-3.1-70B-Instruct 65.08 83.07 62.17 67.72 68.78 67.99
Qwen2.5-32B-Instruct 75.40 84.13 76.72 80.42 79.37 79.10
Qwen2.5-72B-Instruct 76.19 84.66 75.40 78.31 78.31 78.04
HumanEval Llama-3.1-8B-Instruct 63.35 79.88 64.02 74.39 74.39 68.29
Llama-3.1-70B-Instruct 75.61 90.85 76.83 88.42 85.98 84.76
Qwen2.5-32B-Instruct 81.10 93.29 83.54 91.46 90.24 87.80
Qwen2.5-72B-Instruct 82.32 93.90 86.59 90.24 90.85 86.59
BCB Llama-3.1-8B-Instruct 31.67 56.84 34.82 41.84 41.23 39.30
Llama-3.1-70B-Instruct 45.44 62.63 43.86 46.93 47.54 45.88
Qwen2.5-32B-Instruct 45.53 65.18 47.02 49.39 48.95 48.25
Qwen2.5-72B-Instruct 46.67 60.18 47.54 49.04 49.30 48.25
D.6 Additional JETTS results.
In Table 10, we present results full results on JETTS. Concretely, we present (a) a prompt ablation, denoted [CritiquePrompt], where we prompt FARE-20B and gpt-oss-20B for an critique and judgment. We additionally report results for different generators than Llama-3.1-8B-Instruct. Note that unlike FARE-20B, gpt-oss-20B does not natively support prompting without intermediate CoT, making an even comparison with our results presented in Fig. 3 unfeasible. Notably, FARE-20B is the only judge to improve performance over greedy across all generators and benchmarks, regardless of prompt. While FARE-8B is a relatively strong judge, it does not improve generator performance universally, struggling with larger generators on harder benchmarks. The trend of small evaluators struggling in helping larger generators was noted originally in JETTS. Across the board, FARE-20B improves in performance over gpt-oss-20B, sometimes by significant margins (e.g., 49.85 vs 29.93 for Llama-3.1-8B-Instruct MATH performance).
Appendix E Prompts and Examples
E.1 Our evaluation prompts
Below we provide our evaluation prompts for pairwise, step-level, and verification evaluation, along with our direct judgment evaluation prompt for pairwise.
Pairwise evaluation prompt for FARE ### System Prompt Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user prompt displayed below. You will be given assistant A’s answer and assistant B’s answer. Your job is to determine which assistant’s answer is better. If assistant A is better, output [A]. If assistant B is better, output [B]. Here are some rules for evaluation (1) When evaluating the assistants’ answers, identify any mistakes or inaccurate information. Focus on the content each response and select the response that is logically sound and error free. (2) If both responses contain inaccurate information, select the response that arrives at the correct response (3) Avoid any biases, such as order of responses, length, or stylistic elements like formatting Before outputting your final judgment, provide an explanation of your judgment. Your explanation should discuss why your chosen response is better based on the evaluation criteria. The explanation should concretely discuss strengths and weaknesses of both answers. After outputting your explanation, provide your final judgment. Use the following format: Explanation: Your explanation here Verdict: Your final verdict ### User Prompt [User Question] {question} [The Start of Assistant A’s Answer] {response_a} [The End of Assistant A’s Answer] [The Start of Assistant B’s Answer] {response_b} [The End of Assistant B’s Answer]
Direct judgment pairwise evaluation prompt for FARE ### System Prompt Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user prompt displayed below. You will be given assistant A’s answer and assistant B’s answer. Your job is to determine which assistant’s answer is better. If assistant A is better, output [A]. If assistant B is better, output [B]. Here are some rules for evaluation (1) When evaluating the assistants’ answers, identify any mistakes or inaccurate information. Focus on the content each response and select the response that is logically sound and error free. (2) If both responses contain inaccurate information, select the response that arrives at the correct response (3) Avoid any biases, such as order of responses, length, or stylistic elements like formatting Output your final judgment directly. Do not output any explanation or rationale for your decision. Use the following format: Verdict: Your final judgment ### User Prompt [User Question] {question} [The Start of Assistant A’s Answer] {response_a} [The End of Assistant A’s Answer] [The Start of Assistant B’s Answer] {response_b} [The End of Assistant B’s Answer]
Step-level evaluation prompt for FARE ### System Prompt Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user prompt displayed below. You will be given the assistant’s solution to a math problem, which is split into steps, starting with a <step [step number]> tag, where [step number] is indexed from 0. Your job is to identify which step an error occurs, if an error is present. When evaluating the solution, consider each step separately. Evaluate the content of each step for correctness. If you encounter a mistake at <step [step number]>, output [step number] as your Verdict. If the full response is error free, then select step number -1. Avoid any biases, such as length of step, or stylistic elements like formatting. Here are some rules for evaluation. (1) The assistant’s answer does not need to be complete or arrive at a final solution. You may receive a partially complete response. Your job is to assess the quality of each step. (2) When evaluating the assistant’s answer, identify any mistakes or inaccurate information. Focus on the content each step and determine if the step is logically valid. (3) For each step, you should provide an explanation of your assessment. If you find an error, describe the nature and cause of the error. (4) Avoid any biases, such as answer length, or stylistic elements like formatting. Before providing an your final verdict, think through the judging process and output your thoughts as an explanation After providing your explanation, you must output the corresponding step number with an error. Use the following format: Explanation: Your explanation here Verdict: The step number with the error or -1 if no error occurs ### User Prompt [User Question] {question} [The Start of Assistant’s Answer] {response} [The End of Assistant’s Answer]
Reference-based verification evaluation prompt for FARE ### System Prompt Please act as an impartial judge and evaluate if a response provided by an AI assistant (candidate answer) is consistent with a provided reference answer. Your job is to determine is the assistant’s response is consistent with the reference answer. If the response is consistent, output [A]. If the response is incorrect, output [B]. Here are some rules for evaluation. (1) Refer to the given reference answer and determine if the candidate’s answer is consistent with the reference answer. (2) The reference answer is always correct and the question is perfectly valid. Take the reference answer as the ground truth. (3) When determining if the candidate’s answer is consistent with the reference answer, only compare the final answer. Ignore any potential errors in the reasoning processes. (4) Some answers may be expressed in different ways, such as some answers may be a mathematical expression, some answers may be a textual description, as long as the meaning expressed is the same. Before making a judgment, please understand the question and the reference answer first, and then judge whether the candidate’s answer is consistent with the reference answer. (5) Some answers may consist of multiple items, such as multiple-choice questions, multiple-select questions, fill-in-the-blank questions, etc. Regardless of the question type, the final answer will be considered correct as long as it matches the standard answer, regardless of whether the reasoning process is correct. For multiple-select questions and multiple-blank fill-in-the-blank questions, all corresponding options or blanks must be answered correctly and match the standard answer exactly to be deemed correct. Before outputting your final judgment, provide an explanation of your judgment. Your explanation should discuss why the response is correct, incorrect, or invalid. The explanation should concretely discuss reasons for your judgment. After outputting your explanation, provide your final judgment. Use the following format: Explanation: Your explanation here Verdict: Your final judgment of [A] or [B] ### User Prompt <|User Prompt|> {question} <|The Start of Assistant’s Answer|> {response} <|The End of Assistant’s Answer|> <|The Start of Reference Answer|> {reference} <|The End of Reference Answer|>
E.2 Sample evaluation rubric
Here, we provide a sample rubric that was hand-written for SWE-Rank (reddy2025swerank). SWE-Rank data consists of contrastive pairs for training retrieval models. We re-purposed this data into a binary verification task, asking the evaluator if the retrieved code snippet was relevant for editing given a user request. “Positive” samples were assigned “Correct” labels, and “Negative” samples were assigned “Incorrect” labels.
Example hand-written rubric for code retrieval samples Here are some rules for evaluation (1) Determine if the function provided by the assistant is a relevant candidate for editing given the user’s instruction (2) A relevant function is one means that needs to be modified in order to address the issue described in the user’s instruction (3) Modifying a relevant function does not mean is is sufficient to resolve the user’s issue. That is, it is ok if modifying the function does not completely resolve the user issue, but it should make progress towards issue resolution.